40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7730     Entry time: Tue Nov 20 02:57:24 2012
Author: Ayaka, Den, Koji 
Type: Update 
Category: Locking 
Subject: red in arms 

We aligned and locked x and y arms.

MCL loop makes arms lock unstable, adds a lot of noise at frequencies 60-100 Hz. We'll fix it.

At some point we were not able to lock because of ADC overflows of PO signals. They happened if whitening filters were enabled. So we reduced the gain of POX whitening filters down to 36 dB and POY - to 39 dB. Now cavities can be locked with whitening filters.

Also we changed the pedestal of the lens in the beam path to the POX because the beam was too high.

 

arms.png    arms_psd.png

etmxf.png  etmyf.png

itmxf.png     itmyf.png

etmxt.png    etmyt.png

ELOG V3.1.3-